Grade 11Mechanics

Net force acting on particle of mass 2 kg is 10N in north direction.At t=o, particle was moving eastwards with 10m/s. Find displacement and velocity of particle after 2 seconds.

saurabh kharche
13 Years agoGrade 11

2 Answers

naninagaraju sivacharan  lakshmi narsing
13 Years ago

given force =F=10N

mass M=2kg

velocity u=10m/s

then acceleration a=F/M=10/2=5m/s^2

then displacement=s=ut+1/2at^2

here u=initial velocity

t=time

a=acceleration

then we get the displacement is=30m

2.velocity=u+at

then we get velocity as 20m/s.

thanks for asking these type of questions

Imaad
8 Years ago
given force =F=10Nmass M=2kgvelocity u=10im/sthen acceleration a=F/M=10/2=5jm/s^2then displacement=s=ut+1/2at^2here u=initial velocityt=timea=accelerationThen we get displacement =20i+10jAnd velocity(v)=u+at=10i+10jThanks for asking these questions